Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2021-08-04

Observers: Priscila PessiLluís Galbany (night plan)
Support:
TAT:  Nada Ihanec, Emmy Paraskeva, Claudia Gutiérrez, Jacob Teffs,  Michael Fulton,  Janet Chen


EFOSC Calibrations:
Bias, domeflats

Weather:
Start: humidity ~ 29%, wind ~ 7.3m/s, seeing ~ 1.1"
Middle:
 humidity ~ %, wind ~ m/s, seeing ~ "
End: humidity ~ 25%, wind ~ 3.9m/s, seeing ~ 1.5"

Comments (times in UT):
There was a power outage at La Silla which means all systems shut down.
The recovery process took a couple of ours. After opening some time was needed to stabilize telescope temperature.
Once the telescope was stabilized image analysis was performed and observations started.

 UTTarget 
 TypeSetup SlitExposure (s)Airmass
Seeing 
Notes
EFOSC2
        
01:45
LTT7379
std
gr11-13-16
1-1.5
-
1.0
1.1
 Seeing from image quality
 02:03 AT2021ugaclassification
gr13
 1.5900
 1.5 1.8Seeing from image quality
CCSNe
 02:27 AT2021tii classification gr13 1.0 1000 1.5 1.1Seeing from image quality.
Something with hydrogen ?
 02:55 AT2021ugsclassification
 gr131.5
 4001.4
 2.0DIMM started
Noisy spectrum. Trace mounted to the one of the star next to it. I should have considered the awful seeing. Or try a different angle maybe.
 03:15 AT2020usafollowup
 gr131.5
 2 x 2700
 1.6 2.0 Seeing improved dramatically during 2nd exposure
05:02
eg131
std
gr11-13-16
 1-1.5-
 1.21.4
 
05:29
 SN2021tywfollowup
 gr11-16 1 1500 each
 1.5 1.5 
06:32
 SN2020uew followup g-r-i filters
 - - 1.9 1.1 imaging
 08:29 AT2021blzfollowup
 gr131
2 x 2700
 1.5 1.6 
10:10
 AT2021uou classification gr13 11000
 1.11.3
SN Ia
Comments